# A Survey on Mixture of Experts

### v1

@@ 4.2.4. Shared Expert
DeepSpeed-MoE (Rajbhandari et al., 2022) innovatively introduces the Residual-MoE architecture, wherein each token is processed by a fixed expert and another selected through gating, achieving two experts engagement per layer without increasing the communication cost beyond that of top-1 gating. This approach considers the gating-selected MoE expert as an error-correcting adjunct to the fixed dense FFN. A conceptually similar approach, Conditional MoE Routing (CMR), is employed in NLLB (Costa-jussà et al., 2022), which also combines the outputs of dense FFN and MoE layers. This paradigm of integrating fixed FFN with sparse MoE, often referred to as shared expert and illustrated in Figure 5 (b), has gained traction in recent language models such as DeepSeekMoE (Dai et al., 2024), OpenMoE (Xue et al., 2024), Qwen1.5-MoE (Team, 2024b), and MoCLE (Gou et al., 2023), indicating its ascension to a mainstream configuration. Instead of using a single shared expert, DeepSeekMoE (Dai et al., 2024) and Qwen1.5-MoE (Team, 2024b) employ multiple shared experts, due to their fine-grained expert segmentation design.
However, shared expert configuration, while effective in NLP tasks, have not demonstrated the same level of enhancement in vision tasks. Empirical evidence from ScMoE (Cai et al., 2024) indicates that pairing one shared expert with one gating-selected expert yields only comparable performance to standard top-1 MoE in SwinV2-MoE models. Additionally, based on the design of shared expert, ScMoE decouples the MoE process to separately handle the representations from preceding layers and integrate them with the outputs processed by the shared expert of the current layer, thus improving efficiency by facilitating overlap in communication and computation. A comparable method to enhance overlapping is employed in the Dense-MoE hybrid transformer architecture, as delineated in Snowflake Arctic (Team, 2024c), which bears resemblance to the Lora MoE framework discussed in Section 4.3.3 and illustrated in Figure 6 (d).
4.3. Mixture of Parameter-Efficient Experts
@@
LLMs pretrained on generic massive datasets have demonstrated impressive abilities, enabling their deployment across diverse tasks (Gao et al., 2024). However, to tailor a pretrained LLM for a specific downstream task, fine-tuning is essential. Traditional full fine-tuning, which updates all the parameters of the base model, is computationally intensive, especially as model sizes continue to grow (Ding et al., 2022). To address this issue, research into parameter-efficient fine-tuning (PEFT) has emerged, intending to reduce computational demands during the adaptation of a generic pre-trained model to particular tasks (Han et al., 2024). PEFT methods only update a small set of parameters while maintaining the rest of the base model untouched (Lialin et al., 2023). These techniques have achieved state-of-the-art performance across numerous NLP tasks (Liu et al., 2022; Hu et al., 2021).
@@
Despite these successes, PEFT approaches often struggle with generalizing across multiple tasks due to their limited scope of trainable parameters and the potential for catastrophic forgetting (Li et al., 2024b). To mitigate these limitations, a line of mixture of parameter-efficient experts (MoPE) research has emerged, focusing on integrating the MoE framework with PEFT (Li et al., 2024b; Liu et al., 2023). MoPE incorporates the MoE’s gating mechanism and multi-expert architecture, but with each expert constructed using PEFT techniques (Ostapenko et al., 2023). The subtle combination boosts PEFT’s performance under the multi-task scenario (Zhu et al., 2023b). Additionally, by leveraging PEFT for constructing experts, MoPE operates with fewer parameters, achieving greater resource efficiency compared to traditional MoE models (Zadouri et al., 2023).
@@
MoPE harnesses the best of both fields: the task versatility of MoE and the resource efficiency of PEFT (Li et al., 2024b), positioning it as a promising area of study that pushes the boundaries of both fields. In the following subsection, we will give a taxonomy of MoPE, as depicted in Figure 6, based on their placement within the Transformer model architecture. We will then review recent MoPE research, summarizing the methodologies and contributions of each study.
4.3.1. Feed-Forward Network
@@
Following the conventional MoE structure, a series of studies introduce the MoPE framework to the FFN layer of every Transformer block. During the training process, the focus is on optimizing the parameter-efficient experts and the gating mechanism, leaving the rest of the pre-trained model intact. As illustrated in Figure 6(b), the forward process under the MoPE framework integrated with FFN can be expressed as:
@@
One of the pioneering works in this domain, LoRAMoE (Dou et al., 2023b), efficiently applies the MoPE structure to FFN. LoRAMoE integrates a few plug-in LoRA experts into the FFN layer, employing a gating mechanism to orchestrate the experts’ contributions. Realizing the diversity in data distributions, LoRAMoE separates the experts into two distinct groups: one focuses on learning various downstream tasks, and the other is dedicated to aligning pretrained world knowledge with human instructions. To ensure that each group of experts maintains its focus, LoRAMoE defines a localized balancing constraint loss, which preserves the importance of each expert within its group while allowing different groups to concentrate on their respective tasks. This design enables LoRAMoE to effectively resolve the knowledge forgetting issue and enhance model performance on downstream tasks. In a similar vein, AdaMix (Wang et al., 2022) injects a set of Adapter (Houlsby et al., 2019) experts after the FFN layer in each Transformer block. Adapter tuning is a PEFT method that integrates a pair of feed-forward up and down projection matrices into the Transformer block. During fine-tuning, only the incremental Adapter blocks are updated, with the rest of the model unchanged. AdaMix utilizes a stochastic routing policy that randomly selects the projection matrices during training, maintaining computational costs equivalent to a single adapter. To minimize service costs during inference, AdaMix averages the outputs of all experts.
@@
Taking a different approach, MixDA(Diao et al., 2023) includes two training stages to leverage domain-specific knowledge while preserving learned information. During the first stage, MixDA only fine-tunes the domain-adapters that work parallel to the FFN to acquire domain-specific knowledge and keep the world knowledge simultaneously. In the second stage, MixDA introduces a gating network and task-adapters on top of the FFN layer for tailoring the model to specific downstream tasks. This strategy allows for a more nuanced adaptation to the task at hand. LLaVA-MoLE(Chen et al., 2024) extends the application of MoPE to multimodal tasks. It creates a set of LoRA experts for the FFN layer to handle inputs from different domains, enhancing the model’s versatility. LLaVA-MoLE adopts a top-1 routing strategy, activating the most relevant expert based on the router’s output distribution, thus maintaining computational costs close to a standard FFN with LoRA. This framework is effective in addressing data conflicts and consistently surpasses plain-LoRA baselines across diverse data configurations.
@@
Contrasting with the MoPE implementations we have discussed, MixLoRA(Li et al., 2024b) creates a LoRA-MoE framework that closely aligns with the conventional MoE models. Rather than just plugging in multiple lightweight experts, MixLoRA fuses LoRA experts with the shared FFN layer. By leveraging the base weights from a single FFN of the base model, MixLoRA streamlines the creation of the MoPE architecture. Furthermore, MixLoRA implements a high-throughput framework that significantly reduces token computation latency and memory usage during both training and inference, optimizing performance and efficiency.
4.3.2. Attention
@@
Recent studies have demonstrated the effectiveness of extending MoE to the attention layer (Zhang et al., 2022b; Shen et al., 2024, 2023c). Additionally, there is a new line of research has focused on the fusion of MoPE with the attention mechanism to enhance the model’s efficiency and adaptability. For instance, MoELoRA (Luo et al., 2024) applies MoE to the attention mechanism in a resource-efficient manner by leveraging a PEFT method. MoELoRA adopts LoRA (Hu et al., 2021) to construct the experts. LoRA introduces two low-rank matrices to receive incremental updates associated with the task-specific fine-tuning. Only the LoRA matrices are updated while the base model is kept untouched during fine-tuning. Specifically, MoELoRA sets multiple LoRA experts to the Query and Value matrices of the attention mechanism, and utilizes a gating network to activate the top experts related to the specific tasks during both training and inference phases. To alleviate routing randomness, MoELoRA employs a contrastive learning loss to control the training of experts. The contrastive learning loss is designed to accentuate the differences in output distributions between experts, thereby encouraging them to capture diverse features relevant to the downstream tasks. MoELoRA offers a solution for flexibly combining various computational modules tailored to downstream tasks.
Another framework, MoCLE(Gou et al., 2023), aims to resolve task conflicts that arise from the diversity of training tasks of different sources and formats. MoCLE utilizes a clustering model to categorize different tasks and then leverages a router to direct the clustered input to LoRA experts inserted into the Query and Value modules of the attention mechanism. These LoRA experts contain a group of multiple task experts and a universal expert. Each task expert is dedicated to a particular task to reduce task conflicts, while the universal expert, trained on all tasks, helps to maintain model generalization. SiRA(Zhu et al., 2023b) introduces several lightweight LoRA adapters as experts, along with a top gating mechanism. To mitigate load imbalance and over-fitting issues, SiRA incorporates a capacity constraint that limits the number of tokens each expert can process. Additionally, it employs an auxiliary loss to promote load balancing and an expert dropout mechanism to equalize the gating distribution. SiRA provides an efficient and fine-grained approach to improving the quality of LoRA.
4.3.3. Transformer Block
@@
MoV (Zadouri et al., 2023) is one of the notable attempts that combine MoPE with the Transformer block to pursue parameter efficiency. Utilizing the PEFT method, (Liu et al., 2022), MoV introduces tunable vectors that re-scale the Key and Value modules in the attention mechanism, as well as the activation within the FFN. By substituting conventional experts with vectors and updating only these lightweight experts and their corresponding gating during fine-tuning, MoV significantly reduces the computational burden associated with gradient calculations and lessens the memory footprint required for model storage. Similarly, MoLORA (Zadouri et al., 2023) employs multiple LoRA experts to the attention and FFN blocks, outperforming the standard LoRA approach. UniPELT (Mao et al., 2022) proposed a hybrid framework that integrates three representative PEFT methods as experts, namely Adapter (Houlsby et al., 2019), Prefix-tuning (Li and Liang, 2021), and LoRA (Hu et al., 2021). Prefix-tuning is a method that freezes the base model and optimizes the continuous task-specific vectors prepended to the input of the attention. Within the UniPELT framework, LoRA matrices are applied to the weight matrices of Query and Key in the attention mechanism, Prefix vectors are added to the Key and Value modules, and the Adapter block is inserted after the FFN layer. UniPELT leverages different gating mechanisms to dynamically activate the experts, efficiently finding the approaches that best suit the given task.
@@
Further broadening the scope of the LoRA-MoE framework, Omni-SMoLA(Wu et al., 2023a) extends the MoPE with three sets of LoRA experts, each tailored to handle text tokens, visual tokens, and multimodal tokens, respectively. The specialization enables the architecture to enhance performance across various vision-and-language tasks. In the context of MoPE research, the number of experts emerges as a critical hyperparameter influencing downstream task performance (Zadouri et al., 2023; Wu et al., 2024a). Additionally, the use of many experts will lead to redundancy (Chen et al., 2023b). MoLA (Gao et al., 2024) is one of the pioneering work that explores the expert allocation issue. It proposes a LoRA-MoE framework with a Layer-wise Expert Allocation, which enables the flexible employment of varying numbers of experts across different layers. The expert allocation strategy proposed by MoLA further improves the effectiveness of the LoRA-MoE framework. In the specialized field of medical applications, MOELoRA(Liu et al., 2023) tackles the challenges of task variety and high adaptation cost. It integrates LoRA experts and task-motivated gate functions into the attention and FFN of each layer. The gating utilizes task identity to modulate expert contributions, creating unique parameter sets tailored to individual tasks. The design of MOELoRA combines the strengths of both MOE and LoRA, strengthening LLM’s capability in medical domains.
4.3.4. Every Layer
@@
In this context, the approach presented by MoLE(Wu et al., 2024a) provides innovative insights. MoLE identifies that various layers within LoRA exhibit unique features. In response to this finding, MoLE pursues to enhance the composition effect of trained LoRAs by dynamically adjusting the layer-specific weights according to the desired objective. This is achieved by plugging a set of trained LoRAs alongside a gating function into each layer. MoLE treats each layer of trained LoRAs as an individual expert and only trains the gating to learn the optimal composition weights for a specified domain. This dynamic linear composition strategy significantly extends the versatility of LoRA, enabling its application across a broader spectrum of practical scenarios.
4.4. Training & Inference Scheme
@@
Original Training & Inference Scheme. Initial training methodologies, as established in seminal works (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022b; Zoph et al., 2022; Riquelme et al., 2021), involve constructing an MoE model and training it from scratch, with inference directly following the model configurations of training.
@@
The advent of MoE models has introduced novel paradigms for training and inference, enabling a flexible approach that synergizes the strengths of dense and sparse models while mitigating their respective weaknesses. As depicted in Figure 7, we divide the emerging schemes into three distinct categories: Dense-to-Sparse, which entails initiating with dense model training and progressively transitioning to a sparse MoE configuration (Wu et al., 2022; Dua et al., 2022; Komatsuzaki et al., 2022; Tianwen Wei, 2024; Lin et al., 2024; Pan et al., 2024; Nie et al., 2021; Chen et al., 2022d); Sparse-to-Dense, which involves degrading a sparse MoE model to a dense form that is more conducive to hardware implementation for inference (Xue et al., 2022a; Chen et al., 2022b; Huang et al., 2023); and Expert Models Merging, a process of integrating multiple pre-trained dense expert models into a unified MoE model (Li et al., 2022; Sukhbaatar et al., 2024; Wang et al., 2023).
4.4.1. Dense-to-Sparse.

### v2

@@ 4.2.4 Shared Expert
DeepSpeed-MoE [66] innovatively introduces the Residual-MoE architecture, wherein each token is processed by a fixed expert and another selected through gating, achieving two experts engagement per layer without increasing the communication cost beyond that of top-1 gating. This approach considers the gating-selected MoE expert as an error-correcting adjunct to the fixed dense FFN. A conceptually similar approach, Conditional MoE Routing (CMR), is employed in NLLB [78], which also combines the outputs of dense FFN and MoE layers. This paradigm of integrating fixed FFN with sparse MoE, often referred to as shared expert and illustrated in Figure 5 (b), has gained traction in recent language models such as DeepSeekMoE [69], OpenMoE [38], Qwen1.5-MoE [104], and MoCLE [46], indicating its ascension to a mainstream configuration. Instead of using a single shared expert, DeepSeekMoE [69] and Qwen1.5-MoE [104] employ multiple shared experts, due to their fine-grained expert segmentation design. He et al. [111] introduce Partially Dynamic Networks (PAD-Net), iteratively transforming partial parameters of gating-selected experts into static parameters (akin to shared experts) based on their impact on loss values. Zhao et al. [112] introduce HyperMoE, an innovative MoE framework that integrates expert-shared and layer-shared hypernetwork to effectively capture cross-expert and cross-layer information. Additionally, based on the design of shared expert, ScMoE [110] decouples the MoE process to separately handle the representations from preceding layers and integrate them with the outputs processed by the shared expert of the current layer, thus improving efficiency by facilitating overlap in communication and computation. A comparable method to enhance overlapping is employed in the Dense-MoE hybrid transformer architecture, as delineated in Snowflake Arctic [32], which bears resemblance to the LoRA MoE framework discussed in Section 4.3.3 and illustrated in Figure 6 (d).
4.3 Mixture of Parameter-Efficient Experts
@@
LLMs pretrained on generic massive datasets have demonstrated impressive abilities, enabling their deployment across diverse tasks [118]. However, to tailor a pretrained LLM for a specific downstream task, fine-tuning is essential. Traditional full fine-tuning, which updates all the parameters of the base model, is computationally intensive, especially as model sizes continue to grow [185]. To address this issue, research into parameter-efficient fine-tuning (PEFT) has emerged, intending to reduce computational demands during the adaptation of a generic pretrained model to particular tasks [186]. PEFT methods only update a small set of parameters while maintaining the rest of the base model untouched [187]. As an example of PEFT, LoRA [174] introduces two low-rank matrices to receive incremental updates associated with the task-specific fine-tuning. Only the LoRA matrices are updated while the base model is kept untouched during fine-tuning. These techniques have achieved state-of-the-art performance across numerous NLP tasks [188, 174].
@@
Despite these successes, PEFT approaches often struggle with generalizing across multiple tasks due to their limited scope of trainable parameters and the potential for catastrophic forgetting [114]. To mitigate these limitations, a line of mixture of parameter-efficient experts (MoPE) research has emerged, focusing on integrating the MoE framework with PEFT [114, 119]. MoPE incorporates the MoE’s gating mechanism and multi-expert architecture, but with each expert constructed using PEFT techniques [189]. The subtle combination boosts PEFT’s performance under the multi-task scenario [116]. Additionally, by leveraging PEFT for constructing experts, MoPE operates with fewer parameters, achieving greater resource efficiency compared to traditional MoE models [42].
@@
MoPE harnesses the best of both fields: the task versatility of MoE and the resource efficiency of PEFT [114], positioning it as a promising area of study that pushes the boundaries of both fields. In the following subsection, we will give a taxonomy of MoPE, as depicted in Figure 6, based on their placement within the Transformer model architecture. We will then review recent MoPE research, summarizing the methodologies and contributions.
4.3.1 Feed-Forward Network
@@
Following the conventional MoE structure, a series of investigations introduce the MoPE framework to the FFN layer of every Transformer block. During the training process, the focus is on optimizing the parameter-efficient experts and the gating mechanism, leaving the rest of the pretrained model intact. As illustrated in Figure 6(b), the forward process under the MoPE framework integrated with FFN can be expressed as:
@@
One of the pioneering studies in this domain, LoRAMoE [45], efficiently applies the MoPE structure to FFN. LoRAMoE integrates a few plug-in LoRA experts into the FFN layer, employing a gating mechanism to orchestrate the experts’ contributions. Realizing the diversity in data distributions, LoRAMoE separates the experts into two distinct groups: one focuses on learning various downstream tasks, and the other is dedicated to aligning pretrained world knowledge with human instructions. To ensure that each group of experts maintains its focus, LoRAMoE defines a localized balancing constraint loss, which preserves the importance of each expert within its group while allowing different groups to concentrate on their respective tasks. This design enables LoRAMoE to effectively resolve the knowledge forgetting issue and enhance model performance on downstream tasks. In a similar vein, AdaMix [44] injects a set of Adapter [190] experts after the FFN layer in each Transformer block. Adapter tuning is a PEFT method that integrates a pair of feed-forward up and down projection matrices into the Transformer block. During fine-tuning, only the incremental Adapter blocks are updated, with the rest of the model unchanged. AdaMix utilizes a stochastic routing policy that randomly selects the projection matrices during training, maintaining computational costs equivalent to a single adapter. To minimize service costs during inference, AdaMix averages the outputs of all the experts.
@@
Taking a different approach, MixDA[113] includes two training stages to leverage domain-specific knowledge while preserving learned information. During the first stage, MixDA only fine-tunes the domain-adapters that work in parallel with the FFN to acquire domain-specific knowledge while simultaneously retaining the world knowledge. In the second stage, MixDA introduces a gating network and task-adapters on top of the FFN layer for tailoring the model to specific downstream tasks. This strategy allows for a more nuanced adaptation to the task at hand. LLaVA-MoLE[115] extends the application of MoPE to multimodal tasks. It creates a set of LoRA experts for the FFN layer to handle inputs from different domains, enhancing the model’s versatility. LLaVA-MoLE adopts a top-1 routing strategy, activating the most relevant expert based on the router’s output distribution, thus maintaining computational costs close to a standard FFN with LoRA. This framework is effective in addressing data conflicts and consistently surpasses plain-LoRA baselines across diverse data configurations.
@@
Contrasting with the MoPE implementations we have discussed, MixLoRA[114] creates a LoRA-MoE framework that closely aligns with the conventional MoE models. Rather than just plugging in multiple lightweight experts, MixLoRA fuses LoRA experts with the shared FFN layer. By leveraging the base weights from a single FFN of the base model, MixLoRA streamlines the creation of the MoPE architecture. Furthermore, MixLoRA implements a high-throughput framework that significantly reduces token computation latency and memory usage during both training and inference, optimizing performance and efficiency.
4.3.2 Attention
@@
Recent studies have demonstrated the effectiveness of extending MoE to the attention layer [82, 83, 73]. Additionally, there is a new line of research has focused on the fusion of MoPE with the attention mechanism to enhance the model’s efficiency and adaptability. For instance, MoELoRA [47] applies MoE to the attention mechanism in a resource-efficient manner by leveraging a PEFT method, adopting LoRA [174] to construct the experts. Specifically, MoELoRA sets multiple LoRA experts to the Query and Value matrices of the attention mechanism, and utilizes a gating network to activate the top- experts related to the specific tasks during both training and inference phases. To alleviate routing randomness, MoELoRA employs a contrastive learning loss to control the training of experts. The contrastive learning loss is designed to accentuate the differences in output distributions between experts, thereby encouraging them to capture diverse features relevant to the downstream tasks. MoELoRA offers a solution for flexibly combining various computational modules tailored to downstream tasks.
Another framework, MoCLE[46], aims to resolve task conflicts that arise from the diversity of training tasks of different sources and formats. MoCLE utilizes a clustering model to categorize different tasks and then leverages a router to direct the clustered input to LoRA experts inserted into the Query and Value modules of the attention mechanism. These LoRA experts contain a group of multiple task experts and a universal expert. Each task expert is dedicated to a particular task to reduce task conflicts, while the universal expert, trained on all tasks, helps to maintain model generalization. SiRA[116] introduces several lightweight LoRA adapters as experts, along with a top- gating mechanism. To mitigate load imbalance and over-fitting issues, SiRA incorporates a capacity constraint that limits the number of tokens each expert can process. Additionally, it employs an auxiliary loss to promote load balancing and an expert dropout mechanism to equalize the gating distribution. SiRA provides an efficient and fine-grained approach to improving the quality of LoRA.
4.3.3 Transformer Block
@@
MoV [42] is one of the notable attempts that combine MoPE with the Transformer block to pursue parameter efficiency. Utilizing the PEFT method, [188], MoV introduces tunable vectors that re-scale the Key and Value modules in the attention mechanism, as well as the activation within the FFN. By substituting conventional experts with vectors and updating only these lightweight experts and their corresponding gating during fine-tuning, MoV significantly reduces the computational burden associated with gradient calculations and lessens the memory footprint required for model storage. Similarly, MoLORA [42] employs multiple LoRA experts to the attention and FFN blocks, outperforming the standard LoRA approach. UniPELT [117] proposed a hybrid framework that integrates three representative PEFT methods as experts, namely Adapter [190], Prefix-tuning [191], and LoRA [174]. Prefix-tuning is a method that freezes the base model and optimizes the continuous task-specific vectors prepended to the input of the attention. Within the UniPELT framework, LoRA matrices are applied to the weight matrices of Query and Key in the attention mechanism, Prefix vectors are added to the Key and Value modules, and the Adapter block is inserted after the FFN layer. UniPELT leverages different gating mechanisms to dynamically activate the experts, efficiently finding the approaches that best suit the given task.
@@
Further broadening the scope of the LoRA-MoE framework, Omni-SMoLA[43] extends the MoPE with three sets of LoRA experts, each tailored to handle text tokens, visual tokens, and multimodal tokens, respectively. The specialization enables the architecture to enhance performance across various vision-and-language tasks. In the context of MoPE research, the number of experts emerges as a critical hyperparameter influencing downstream task performance [42, 48]. Additionally, the use of many experts may lead to redundancy [192]. MoLA [118] is one of the pioneering work that explores the expert allocation issue. It proposes a LoRA-MoE framework with a Layer-wise Expert Allocation, which enables the flexible employment of varying numbers of experts across different layers. The expert allocation strategy proposed by MoLA further improves the effectiveness of the LoRA-MoE framework. In the specialized field of medical applications, MOELoRA[119] tackles the challenges of task variety and high adaptation cost. It integrates LoRA experts and task-motivated gate functions into the attention and FFN of each layer. The gating utilizes task identity to modulate expert contributions, creating unique parameter sets tailored to individual tasks. The design of MOELoRA combines the strengths of both MoE and LoRA, strengthening LLM’s capability in medical domains. Liu et al. [120] design a novel framework, named Intuition-MoR1E, which leverages the inherent semantic clustering of instances to emulate cognitive processes in the human brain for multitasking purposes. This framework provides implicit guidance to the gating mechanism, thereby enhancing feature allocation. Furthermore, they introduce a cutting-edge rank-1 expert architecture. This architecture advances beyond the conventional 1-1 mapping of two weight matrices and in LoRA expert composition, facilitating a flexible combination of any with any to form an expert. They implement MoE in the transformer blocks, specifically targeting the Query, Key, Value, and FFN modules. Xu et al. [121] present Multiple-Tasks Embedded LoRA (MeteoRA), a scalable and efficient framework that embeds multiple task-specific LoRA adapters into the base LLM using a full-mode MoE architecture. This framework incorporates custom GPU kernel operators to accelerate LoRA computation while maintaining memory overhead.
4.3.4 Every Layer
@@
In this context, the approach presented by MoLE[48] provides innovative insights. MoLE identifies that various layers within LoRA exhibit unique features. In response to this finding, MoLE pursues to enhance the composition effect of trained LoRAs by dynamically adjusting the layer-specific weights according to the desired objective. This is achieved by integrating a set of trained LoRAs and a gating function into each layer. MoLE treats each layer of trained LoRAs as an individual expert and only trains the gating to learn the optimal composition weights for a specified domain. This dynamic linear composition strategy significantly extends the versatility of LoRA, enabling its application across a broader spectrum of practical scenarios.
4.4 Training & Inference Scheme
@@
Original Training & Inference Scheme. Initial training methodologies, as established in seminal works [27, 28, 36, 37, 8], involve constructing an MoE model and training it from scratch, with inference directly following the model configurations of training.
@@
The advent of MoE models has introduced novel paradigms for training and inference, enabling a flexible approach that synergizes the strengths of dense and sparse models while mitigating their respective weaknesses. As depicted in Figure 7, we divide the emerging schemes into three distinct categories: Dense-to-Sparse, which entails initiating with dense model training and progressively transitioning to a sparse MoE configuration [122, 123, 49, 67, 106, 64, 62, 124]; Sparse-to-Dense, which involves degrading a sparse MoE model to a dense form that is more conducive to hardware implementation for inference [52, 53, 127]; and Expert Models Merging, a process of integrating multiple pretrained dense expert models into a unified MoE model [129, 54, 130].
4.4.1 Dense-to-Sparse